undeadknight26 (undeadknight26):

Barbara is a waitress and she usually takes between 5 and 9 tables at a time. She found that when she is assigned five tables, she averages $15 in tips per table per shift. If she is assigned more tables, the amount of attention she gives to each table decreases. Therefore, she averages less in tips per table. Suppose each additional table causes every table to bring in $3 less in tips. Which of the following equations could be used to find the amount of money she would make in tips if she takes 5 + x tables? T(x) = -3x2 + 75 T(x) = x2 + 5x T(x) = 20 - x T(x) = 3x2 + 75

undeadknight26 (undeadknight26):

@Zale101

undeadknight26 (undeadknight26):

I thought this was A.

Miracrown (miracrown):

You're right! It is A :)

undeadknight26 (undeadknight26):

Can you help with a few more @Miracrown ?

Miracrown (miracrown):

Sure, lets see ...

undeadknight26 (undeadknight26):

Luke had money invested into two separate companies in the stock market. At the beginning of the week, both companies had the same amount of money in their accounts. Then, each stock began to decrease daily. Company A's stock decreased by a dollars a day and Company B's stock decreased by b dollars a day. After x days, which of the following expressions represents how much more money Company B lost than Company A? bx - a b - ax (a - b)x (b - a)x

Miracrown (miracrown):

What are your thoughts

undeadknight26 (undeadknight26):

Wasnt really sure where to start on this one...But if i were to think anything it would be C.

Miracrown (miracrown):

hmmm.. it seems that could be a different one but very similar?

undeadknight26 (undeadknight26):

Oh so D?

Miracrown (miracrown):

Yep

undeadknight26 (undeadknight26):

Which expression represents the following sequence of operations on a number ? Subtract 2 from then divide the answer by 7.

undeadknight26 (undeadknight26):

I think (x - 2)/7?

Miracrown (miracrown):

Subtract two from what?

undeadknight26 (undeadknight26):

subtract 2 from x

Miracrown (miracrown):

Ok, then it is correct!

undeadknight26 (undeadknight26):

Just like 3 more?

Miracrown (miracrown):

Okie

undeadknight26 (undeadknight26):

At the movie theater, the price of an adult ticket is "a" dollars and the price of child's ticket is "b" dollars. If x adults and y children go to the movies, which of the following expressions represents the total number of adults and children who go to the movies? x + a ax + by a + b x + y

undeadknight26 (undeadknight26):

I thought D

Miracrown (miracrown):

Correct

undeadknight26 (undeadknight26):

If a and b are integers, which of the following expressions represents the sum of two odd integers? 2(a + b) + 1 2(a + b + 1) a + b + 2 a + b I thought D again.

Miracrown (miracrown):

Not quite, while that is sometimes the case it is not always the case.

undeadknight26 (undeadknight26):

Then B maybe?

Miracrown (miracrown):

Hint: among your answers there is an expression guaranteed to be odd no matter what.

Miracrown (miracrown):

Try again there are two ways you can solve this problem

Miracrown (miracrown):

One is to consider different values of a and b, like both even, both odd, one even and the other odd. Another is to remember that 2*(any integer) +1 is always odd.

undeadknight26 (undeadknight26):

Ohhh so its A ._.

Miracrown (miracrown):

Yerp .-.

undeadknight26 (undeadknight26):

It was b ._.

Miracrown (miracrown):

Since a and b are integers , there sum is an integer , since 2*integer +1 is always odd a is the answer.

Miracrown (miracrown):

A way to remember that formula is : 2*anything is even , and the next number in sequence +1 is always odd.

Miracrown (miracrown):

It can't be b that is guaranteed to be even. 2*(anything ) by definition is even.

undeadknight26 (undeadknight26):

Well the test said it was B so this test is flawed. The length of a rectangle is 4 units less than 7 times its width. The area of the rectangle is 188 square units. Which of the following equations can be used to find w, the width of the rectangle? 8w - 28 = 188 8w - 4 = 188 7w2 - 4w = 188 7w2 - 28w = 188

undeadknight26 (undeadknight26):

Last one i think its B again.

Miracrown (miracrown):

Wait I may have misread the problem. Yes , sorry I misread the problem. I thought it was asking which one of those must be odd. If a and b are integers.

Miracrown (miracrown):

I will explain b

Miracrown (miracrown):

We need to express the sum of two odd numbers using a and b. One such number is : 2a+1

undeadknight26 (undeadknight26):

Ohhhh i see...

Miracrown (miracrown):

And you do realize that we're not allowed to help you with a test. It is designed for you, not for someone else. But I can help confirm your answers :)

undeadknight26 (undeadknight26):

lol exactly why im giving you what i think... Dont worry ive been around OS a while lol

Miracrown (miracrown):

I know dood :^)

undeadknight26 (undeadknight26):

Now that im looking at the last one i asked im thinking its either C or B...Im so confused...

Miracrown (miracrown):

Let me check ...

Miracrown (miracrown):

B should be auto eliminated. This is because we require a quadratic expression in terms of w. A = l*w , even when you get l in terms of w , when plugging back in you will get a quadratic in w.

undeadknight26 (undeadknight26):

O.O Well thats something i should def make a note of.

Miracrown (miracrown):

Yerp. So What did you come up with?

undeadknight26 (undeadknight26):

It cant be C or D since thats a hyperbola? So it would have to be A?

Miracrown (miracrown):

Nope, think again ---- it has to be C or D , those are the only ones with a w^2 in them.

undeadknight26 (undeadknight26):

Im lost now ._.

Miracrown (miracrown):

let's do this formally instead of relying on answer choices

undeadknight26 (undeadknight26):

Roger

undeadknight26 (undeadknight26):

Done and the answer is C :) Thanks for your help mate